OpenStudy (anonymous):

A triangle has an area of 20 cm2. The height of the triangle is three cm longer than the base. Find the height and base of the triangle.

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the general formula for the area of a triangle?

OpenStudy (anonymous):

A=hbb over 2

jimthompson5910 (jim_thompson5910):

it should be A = (b*h)/2

jimthompson5910 (jim_thompson5910):

"The height of the triangle is three cm longer than the base" So, height = base + 3 h = b+3

jimthompson5910 (jim_thompson5910):

A = (b*h)/2 A = (b*(b+3))/2 20 = (b*(b+3))/2 20*2 = b*(b+3) 40 = b*(b+3) and I'll let you finish up

OpenStudy (anonymous):

we dont know the base ?

jimthompson5910 (jim_thompson5910):

you solve that equation to get the base

OpenStudy (anonymous):

divide

jimthompson5910 (jim_thompson5910):

40 = b*(b+3) 40 = b^2 + 3b 0 = b^2 + 3b - 40 b^2 + 3b - 40 = 0 Now factor and solve

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b= -8,5???

jimthompson5910 (jim_thompson5910):

toss out the negative solution, you can't have a negative base

jimthompson5910 (jim_thompson5910):

so b = 5 is the only practical solution

jimthompson5910 (jim_thompson5910):

h = b+3 h = 5+3 h = 8

jimthompson5910 (jim_thompson5910):

b = 5 h = 8

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome